OpenStudy (spring98):

HELP PPL!! ----> The length of a rectangle is 7 units and its width is 5 units. What is the approximate length of the diagonal of the rectangle? 4.9 units 6 units 8.6 units 9 units

OpenStudy (spring98):

@Agl202

OpenStudy (anonymous):

So what you want to use here is Pythagoras' Theorem, because you're looking at a right angled triangle.  Have you heard f Pythagoras' Theorem?

OpenStudy (spring98):

yes but i don't know how to use it

OpenStudy (anonymous):

so in a rectangle you have a length of 7 and a width of 5 so your long side would by 7 units long and your short side would be 4 units long, correct?

OpenStudy (anonymous):

Okay, so the equation is:\[c ^{2} = a ^{2} + b ^{2}\] where c is the diagonal and a and b are the other sides

OpenStudy (anonymous):



OpenStudy (anonymous):

So in this case the diagonal^2 = 7^2 + 5^2

OpenStudy (anonymous):

@Spring98

OpenStudy (spring98):

oh sorry

OpenStudy (spring98):

so is it 4.9 units

OpenStudy (spring98):

@Ziploxk plz is it 4.9 units?

OpenStudy (anonymous):

i got ~8.6 @Spring98

OpenStudy (anonymous):

\[c^2=a^2+b^2\]

OpenStudy (spring98):

oh oops i did it wrong k i understand now

OpenStudy (anonymous):

then you would square root c^2 to get c